# POV Learning: Individual Alignment of Multimodal Models using Human Perception

### v1

@@ 9 Future work
The most obvious direction for future work is towards generalization of the approaches presented in this paper to other tasks and models. In a first direction, a reliable and cheaper (in comparison to conducting an eye tracking study) method for capturing human fixation sequences or other cues for individual alignment has to be developed. Without that, the benefits of perception-guided POV models cannot be tested in settings where multimodal stimuli are observed “in the wild”. More generally, identifying and obtaining valuable alignment signals is a future challenge for further research in the field of individual alignment and POV learning.
Another interesting future study should be concerned with the performance of GPT4. Apparently, the proposed task and few-shot prompts did not result in a convincing performance. Additional approaches to individualize MLLM’s inferences via fine-tuning or persona-prompts need to be investigated.

### v2

@@ 9 Future work
The most obvious direction for future work in POV learning is towards the generalization of the presented approaches to other forms of individual alignment signals. In a first step, a reliable and less costly method (in comparison to conducting an eye tracking study) for capturing individual human alignment has to be investigated. Ultimately, the benefits of perception-guided POV models must be tested in settings where multimodal stimuli are observed “in the wild”. Identifying and obtaining valuable alignment signals is a future challenge for further research in the field of individual alignment and POV learning.
Another promising future study could be concerned with improving the performance of Multimodal LLMs like GPT4. So far, the proposed task and the few-shot prompts did not result in a convincing performance. Additional approaches to aligning MLLMs’ inferences to individual preferences via post-training need to be investigated.